Find the dimensions of a rectangle if its area is 72m squared and it's perimeter is 34m
Answer by amoresroy(361)

Let x = the width of the rectangle
y = the length of the rectangle
Equations:
x*y = 72
2x +2y = 34 or x + y = 17
--------------------
Substitute y = 72/x in 2nd equation to get
x + 72/x = 17
Multiply by x to get
x^2 + 72 = 17x
x^2 -17X + 72 = 0
Solve by factoring
(x-9)(x-8) = 0
x = 8 meters
y = 9 meters